OpenStudy (anonymous):

3/4+x=7/8 does anyone know how to solve this equation?

OpenStudy (jdoe0001):

multiply both sides by "8", and see what you get :/

OpenStudy (anonymous):

so i multiply 3/4 and 7/8 by 1/8?

OpenStudy (jdoe0001):

no, just plain ol` 8 :)

OpenStudy (nurali):

\[\frac{ 3 }{ 4+x }=\frac{ 7 }{ 8 }\] Cross multiply \[8*3=7(4+x)\] \[24=28+7x\] \[7x=-4\] \[x=\frac{ -4 }{ 7 }\]

OpenStudy (anonymous):

GOOD